A Bibliography of Supercomputing ’93

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

04 September 2023 Version 1.05

Abstract

This bibliography records articles presented at the Supercomputing ’93 conference.

Title word cross-reference

N [WS93].

*T [PBGB93].

-Body [WS93].

1 [SJGH93]. 1024 [LM93]. 1024-Node [LM93].

2 [KT93].

3 [Van93, WOU^{+}93]. 3-D [WOU^{+}93]. 3-Dimensional [Van93]. 3D [CMBC93].

5 [KTR93, LTGJB93, LM93].

77 [SWBO93].

860 [Van93].

90D [BCF^{+}93]. 90D/HPF [BCF^{+}93]. ’93 [IEE93, Sha93].

Access [CCHT93, CFPB93]. Acess [cC93]. Adaptive [DFWM93, NX93, SHHG93, SJ93]. ADI [Van93]. Advanced [BW93]. algebra [DPW93]. Algorithm [BPS93, TSBP93, WS93]. Algorithms [KHJ^{+}93, Lv93]. Alignment [CGS93]. all-port [RJMC93]. Allocator [NP93]. Analysis [HS93, NSS93, PP93, SGC93]. Analytical [CQ93]. Application [JBBH93, KT93, RKWF93]. Applications [ASS93, GGL93, PP93, PSM93, PY93].

Applications-Driven [GGL93]. Approach [RG93]. Arbitrary [cC93]. Arbitrary-View-Projection [cC93]. Architectures [BKP93, DGL^{+}93, PY93, iMSG^{+}93]. Array [MWBL93]. Arrays [CGS93]. Assessing [TGJ93]. ASURA [iMSG^{+}93]. Automatic [Dag93].

Balancing [DFWM93]. Bandwidth [BR93, Web93]. Barrier [Gra93]. Barrier-breaking [Gra93]. Based [PSM93]. be [TGJ93]. Bell [HKS93]. benchmark [SGC93]. Bispectrum [MWBL93]. Block [ASS93, RG93]. Block-Oriented [RG93]. blocks [PBGB93]. Body [WS93]. Boltzmann [LM93]. breaking [Gra93]. building [PBGB93].

C [BBG^{+}93, Sta93, CDG^{+}93]. C/C [Sta93]. C90 [LCL^{+}93]. C916 [Gra93]. Cache [Bal93, DK93a, TGJ93, iMSG^{+}93]. calculations [LCL^{+}93]. Carlo [Din93, SMBZ93]. Challenge [BK93, WOU^{+}93]. Characteristics [PP93]. Chemical [GL93]. Cholesky [RG93]. CM [KTR93, KT93, LM93]. CM-2 [KT93]. CM-5 [KTR93, LM93]. Coherence [Bal93, DK93a]. collective [RJMC93]. Common [FRS^{+}93]. Communication [KTR93, PSC93]. Comparison [SJGH93]. Compilation [PSC93]. Compile [TGJ93]. Compile-Time [TGJ93]. Compiler [ASS93, BW93, BCF^{+}93, HKT93, FRS^{+}93]. Components [Dag93]. Computation [EMO^{+}93, KTR93]. Computational [PSM93, Set93]. Computations [BW93]. Compute [Mil93]. Computer [Was93]. Computers [BCF^{+}93, Bro93, NXK93]. Computing [Bai93, BK93, CLF^{+}93, LS93a, Che93, PBGB93, RKWF93]. Concert [KC93]. Concert-Efficient [KC93]. concurrent [KC93]. Conflicts [TGJ93]. Conjugate [Lv93]. Connected [Dag93].

Connection [LTGJB93]. Considerations [Web93]. consortium [FRS^{+}93]. Constrained [BR93]. Convergence [Sta93]. Copy [TGJ93]. Copying [TGJ93]. Corresponding [SMBZ93]. Cost [SSL^{+}93]. CPU [LCL^{+}93]. CRAY [Gra93, LCL^{+}93]. current [BDG^{+}93].

D [HKT93, WOU^{+}93]. DASH [SJGH93]. Data [CMMZ93, CGS93, FK93, OHA93, PSC93, TGJ93, TSBP93, RJMC93]. Data-Parallel [CGS93, TSBP93]. Datasets [CCHT93]. Davidson [SF93]. Daylighting [SMBZ93]. Debuggers [Bro93]. Decentralized [LML^{+}93]. Dedicated [LS93a]. Delaunay [TSBP93]. Design [BdC93, BCF^{+}93, NXK93, SMBZ93, DPW93]. Designing [RR93]. Development [LML^{+}93]. Dimensional [TSBP93, Van93]. direction [BDG^{+}93]. Disk [LLS^{+}93]. Distributed [BCF^{+}93, EMO^{+}93, LS93a, Lv93, LS93b, ZI93, Che93, iMSG^{+}93]. Distributed-Memory [LS93b]. distribution [RJMC93]. Distributions [CMMZ93]. Divergence [Sta93]. Do [BR93]. Driven [GGL93, PY93]. Dynamic [CMMZ93, DFWM93]. Dynamics [LTGJB93, WOU^{+}93].

Effect [Was93]. Efficient [KC93, KHJ^{+}93, MY93, RJMC93, RG93, Van93]. Eigenvalue [SF93]. Electromagnetic [CLF^{+}93]. Element [DFWM93]. Eliminate [TGJ93]. Empirical [SJGH93]. Enhanced [PM93]. Envelope [BPS93]. Environment [CLF^{+}93, JBBH93, PSM93, RAA^{+}93, SBS^{+}93]. Environments [NSS93]. Equation [LM93]. evaluation [BdC93]. Execution [PY93, SWBO93]. Execution-Driven [PY93]. Experiences [HKT93, BDG^{+}93]. Experiments [KT93]. extensions [DPW93]. External [LLS^{+}93].

Factorization [RG93]. Fast [OHA93, SHHG93]. Feasibility [LS93a]. Feed [MY93]. Feed-Forward [MY93]. FFT [Mil93]. Fields [CMBC93]. File [CFPB93]. Files [CFPB93]. Finite [DFWM93]. Fluid [Set93, LCL^{+}93]. Fortran [BKP93, BCF^{+}93, CMMZ93, CKZ93, HKT93, SWBO93]. Fortran-S [BKP93]. Forward [MY93]. fully [SJ93]. Future [DGL^{+}93, BDG^{+}93].

GFlops [LTGJB93]. Gigabit [KT93]. Gigaflops [LM93]. Global [NP93, Was93, ZI93]. Gordon [HKS93]. Gradient [Lv93]. Grand [WOU^{+}93]. Graphics [STB^{+}93]. Grids [Dag93].

Hardware [PM93, KC93]. Hashed [WS93]. Heterogeneous [Che93, RKWF93]. Hetrogeneous [BK93]. High [Bro93, CLF^{+}93, CHW93, DPW93, FRS^{+}93, CKZ93]. HIPPI [STB^{+}93]. HNC [MWBL93]. Homogeneous [LS93a]. Hot [HS93]. HPF [BCF^{+}93]. hypercubes [RJMC93].

I/O [BdC93, GGL93, PP93]. Image [Web93]. Implementation [BCF^{+}93, TSBP93, Van93]. Implementing [BBG^{+}93]. Implementor [CKZ93]. incremental [NX93]. industrial [Gra93]. Information [Sha93]. Infrastructure [Sha93]. Integrated [PBGB93]. Integrating [FK93]. Intensity [Mil93]. Interactive [CLF^{+}93]. Interconnection [RR93]. Interface [BKP93, STB^{+}93, Ano93]. iPSC [Van93]. iPSC/860 [Van93]. Issues [NXK93, RKWF93].

Job [NSS93].

Kendall [SJGH93]. Knowledge [DK93a]. KSR [SJGH93]. KSR-1 [SJGH93]. KSR1 [SGC93]. languages [FRS^{+}93, KC93]. LAPACK [DPW93]. Large [CCHT93, DK93b, HS93, KHJ^{+}93, SF93, WOU^{+}93]. Large-Scale [WOU^{+}93]. Latency [Web93]. Lectures [HKS93]. Level [RR93]. Libraries [DK93b]. Library [NXK93]. linear [DPW93]. LLNL [CMBC93]. Load [DFWM93]. Local [DK93a]. logging [NX93]. Low [SSL^{+}93]. Low-Cost [SSL^{+}93].

Machine [LTGJB93]. Machines [LS93b]. Massively [BBFH93, Bal93, DFWM93, Din93, NXK93, Set93]. Massively-Parallel [BBFH93]. Matrices [BPS93, KHJ^{+}93]. Matrix [Lv93]. Matrix-Vector [Lv93]. Matter [WOU^{+}93]. Mechanics [PSM93, Set93]. Mechanism [cC93]. Memory [BKP93, BCF^{+}93, cC93, HS93, Lv93, LS93b, ZI93, iMSG^{+}93]. meshes [SJ93]. Message [Ano93, NX93]. message-passing [NX93]. Method [DFWM93, SHHG93, SF93, Van93]. Micro [SGC93]. microprocessors [Bai93]. MIMD [BCF^{+}93]. minisymposium [Che93]. Mobile [CGS93]. Model [SWBO93]. Modeling [Was93]. Models [EMO^{+}93]. Molecular [LTGJB93]. Monitoring [RAA^{+}93]. Monte [Din93, SMBZ93]. MP [LCL^{+}93]. MPI [Ano93]. MPPs [LLS^{+}93]. Much [BR93]. Multi [LCL^{+}93, RR93]. Multi-CPU [LCL^{+}93]. Multi-Level [RR93]. Multicomputers [CQ93]. Multiplication [Lv93]. Multipole [SHHG93]. multiprocessor [iMSG^{+}93]. Multiprocessors [HS93, SJGH93].

National [Sha93]. Need [BR93]. Network [BR93]. Networks [RR93]. Neuroscience [KT93]. NII [Sha93]. Node [LM93]. Non [LS93a]. Non-Dedicated [LS93a]. Novel [cC93]. November [IEE93]. Numerical [MWBL93].

O [BdC93, GGL93, PP93]. object [DPW93, KC93]. object-oriented [DPW93, KC93]. Oct [WS93]. Oct-Tree [WS93]. Optimal [LML^{+}93, SJ93]. Optimizations [BW93]. Oregon [IEE93]. Oriented [RG93, DPW93, KC93]. overview [DPW93].

Packaging [RR93]. Panel [CCHT93]. Paragon [RAA^{+}93]. Parallel [BBFH93, Bal93, BBG^{+}93, CGS93, CFPB93, CDG^{+}93, DFWM93, Din93, DGL^{+}93, GGL93, LML^{+}93, MY93, NSS93, NXK93, OHA93, PY93, RG93, SWBO93, Set93, SHHG93, SF93, TSBP93, WS93, Web93, BdC93, FRS^{+}93, PBGB93, BBG^{+}93, Sta93]. Parallelism [FK93]. Partitioning [Dag93, PSC93, ZI93]. Passing [Ano93, NX93]. Performance [BR93, BCF^{+}93, Bro93, CLF^{+}93, CKZ93, CQ93, CHW93, KTR93, NSS93, RAA^{+}93, DPW93, FRS^{+}93, Gra93]. Perspective [BBFH93]. PIPADS [SSL^{+}93]. plasma [LCL^{+}93]. Policies [NSS93]. Polygon [OHA93]. port [RJMC93]. Portland [IEE93]. Power [LML^{+}93]. Practical [LLS^{+}93]. Prediction [CQ93, SWBO93]. Preliminary [HKT93]. Pricing [LML^{+}93]. primitives [BdC93]. Prize [HKS93]. Problem [SF93]. problems [Gra93]. Proceedings [IEE93]. Processing [MWBL93, Web93]. Processor [MWBL93]. Processors [Bal93, MY93, Set93]. Production [PP93]. Program [LML^{+}93]. Programmable [STB^{+}93]. Programming [CDG^{+}93, KC93]. Programs [CGS93, NX93]. Project [WOU^{+}93]. Projection [cC93]. Propagation [EMO^{+}93]. PVM [BDG^{+}93, EMO^{+}93].

Quantum [Din93, WOU^{+}93].

Real [SSL^{+}93]. Real-Time [SSL^{+}93]. Reducing [SF93]. Reduction [BPS93]. Register [NP93]. Relativistic [WOU^{+}93].

Remote [CLF^{+}93]. Rendering [cC93, CMBC93, MY93, OHA93]. replay [NX93]. Replicated [CGS93]. Research [SJGH93]. Results [BCF^{+}93]. Reuse [PSC93]. RISC [Bai93]. Robotics [Web93]. routed [RJMC93]. routing [SJ93]. Runtime [ASS93, BBG^{+}93, PSC93, FRS^{+}93, KC93].

S [BKP93]. Scalable [BBG^{+}93, NXK93]. Scalar [CMBC93]. Scale [HS93, WOU^{+}93]. Scattering [CLF^{+}93]. Schedule [PM93, PSC93]. Scheduler [NP93]. Scheduler-Sensitive [NP93]. Scheduling [LS93b, NSS93]. Scheme [Bal93]. Scientific [Bai93, CCHT93, GL93, PP93, SMBZ93]. Self [LS93b]. Self-Scheduling [LS93b]. Sensitive [NP93]. Shared [BKP93, HS93, LLS^{+}93, iMSG^{+}93]. Should [TGJ93]. Signal [MWBL93]. SIMD [MWBL93]. Simulation [CLF^{+}93, PY93]. Simulations [Din93]. Smart [CCHT93]. SNAP [MWBL93]. Software [RKWF93]. Solutions [WOU^{+}93]. Solving [LM93]. Sort [LLS^{+}93]. Space [ZI93]. Sparse [BPS93, BW93, RG93, SF93]. Spectral [BPS93]. Split [CDG^{+}93]. Split-C [CDG^{+}93]. Spot [HS93]. Square [SJGH93]. Stanford [SJGH93]. Static [PP93]. status [BDG^{+}93]. Steering [JBBH93]. stock [KC93]. Storage [CHW93]. Striping [DK93b]. Structure [WOU^{+}93]. Structured [ASS93]. Suitable [Bal93]. Summary [Che93]. Supercomputer [STB^{+}93]. Supercomputers [BBFH93, Din93]. Supercomputing [IEE93, KM93, NSS93, PSM93, SBS^{+}93, Sha93]. Supercomputing-Based [PSM93]. Superscalar [PM93]. Support [ASS93, FRS^{+}93, KC93]. Synchronization [SF93]. System [BBG^{+}93, CLF^{+}93, CFPB93, CHW93, LS93a]. Systems [BBG^{+}93, Din93, GL93, ZI93, RKWF93].

Table [PM93]. Tape [DK93b]. Task [FK93]. Technique [TGJ93]. Techniques [MY93, PSC93]. Three [TSBP93]. Three-Dimensional [TSBP93]. Time [SSL^{+}93, TGJ93]. Tool [SSL^{+}93]. Tools [PY93]. Transposition [KHJ^{+}93]. Tree [WS93]. Triangulation [TSBP93]. turbulence [LCL^{+}93].

Unstructured [Dag93]. Used [TGJ93]. User [BBFH93]. Users [CKZ93]. Uses [SBS^{+}93]. Using [DK93a, EMO^{+}93].

VASE [JBBH93]. Vector [BBFH93, CMBC93, Lv93, MY93]. Vesta [CFPB93]. Via [SMBZ93]. Videoconferencing [SBS^{+}93]. Vienna [CMMZ93]. View [cC93]. Virtual [BKP93]. Visualization [CLF^{+}93, GL93, JBBH93, PSM93, SMBZ93, SSL^{+}93]. Volume [cC93, CMBC93, MY93]. vs [BBFH93].

Warming [Was93]. Wave [EMO^{+}93]. Weather [SWBO93]. Worklaod [PP93]. Workshop [CKZ93]. World [KM93]. wormhole [RJMC93, SJ93]. wormhole-routed [RJMC93].

Y-MP [LCL^{+}93].

References

Anonymous:1993:MMP

[Ano93] Anonymous. MPI: A Message Passing Interface. In IEEE [IEE93], pages 878–883. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Agrawal:1993:CRS

[ASS93] Gagan Agrawal, Alan Sussman, and Joel Saltz. Compiler and runtime support for structured and block structured applications. In IEEE [IEE93], pages 578–587. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Bailey:1993:RMS

[Bai93] David H. Bailey. RISC microprocessors and scientific computing. In IEEE [IEE93], pages 645–654. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Baldwin:1993:CCS

[Bal93] Reid Baldwin. A cache coherence scheme suitable for massively parallel processors. In IEEE [IEE93], pages 730–739. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Bailey:1993:MVP

[BBFH93] D. H. Bailey, E. D. Brooks, D. W. Forslund, and R. J. Harrison. Massively-parallel vs. parallel vector supercomputers: A user’s perspective. In IEEE [IEE93], pages 918–920. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Bodin:1993:IPC

[BBG^{+}93] F. Bodin, P. Beckman, D. Gannon, S. Yang, S. Kesavan, A. Malony, and B. Mohr. Implementing a Parallel C++ runtime system for scalable parallel systems. In IEEE [IEE93], pages 588–597. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Bozkus:1993:FHC

[BCF^{+}93] Zeki Bozkus, Alok Choudhary, Geoffrey Fox, Tomasz Haupt, and Sanjay Ranka. Fortran 90D/HPF compiler for distributed memory MIMD computers: Design, implementation, and performance results. In IEEE [IEE93], pages 351–360. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Bordawekar:1993:DEP

[BdC93] Rajesh Bordawekar, Juan Miguel del Rosario, and Alok Choudhary. Design and evaluation of primitives for parallel I/O. In IEEE [IEE93], pages 452–461. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Beguelin:1993:PEC

[BDG^{+}93] A. Beguelin, J. Dongarra, A. Geist, R. Manchek, S. Otto, and J. Walpole. PVM: Experiences, current status and future direction. In IEEE [IEE93], pages 765–766. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Berman:1993:HCC

[BK93] F. Berman and T. Kitchens. The hetrogeneous computing challenge. In IEEE [IEE93], pages 616–623. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Bodin:1993:FFI

[BKP93] F. Bodin, L. Kervella, and T. Priol. Fortran-S: A Fortran interface for shared virtual memory architectures. In IEEE [IEE93], pages 274–283. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Barnard:1993:SAE

[BPS93] Stephen T. Barnard, Alex Pothen, and Horst D. Simon. A spectral algorithm for envelope reduction of sparse matrices. In IEEE [IEE93], pages 493–502. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Boothe:1993:PBC

[BR93] B. Boothe and A. Ranade. Performance on a bandwidth constrained network: How much bandwidth do we need? In IEEE [IEE93], pages 906–915. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Brown:1993:DHP

[Bro93] Jeffrey S. Brown. Debuggers for high performance computers. In IEEE [IEE93], pages 442–445. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Bik:1993:ACO

[BW93] Aart J. C. Bik and Harry A. G.

Wijshoff. Advanced compiler optimizations for sparse computations. In IEEE [IEE93], pages 430–439. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Chiueh:1993:NMA

[cC93] Tzi cker Chiueh. A novel memory acess mechanism for arbitraryview-projection volume rendering. In IEEE [IEE93], pages 690–698. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Christian:1993:PSA

[CCHT93] Carol Christian, Robert Cromp, Carol Hunter, and Lloyd Treinish. Panel on smart access to large scientific datasets. In IEEE [IEE93], pages 372–373. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Culler:1993:PPS

[CDG^{+}93] David E. Culler, Andrea Dusseau, Seth Copen Goldstein, Arvind Krishnamurthy, Steven Lumetta, Thorsten von Eicken, and Katherine Yelick. Parallel programming in Split-C. In IEEE [IEE93], pages 262–273. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Corbett:1993:PAF

[CFPB93] Peter F. Corbett, Dror G. Feitelson, Jean-Pierre Prost, and Sandra Johnson Baylor. Parallel access to files in the Vesta file system. In IEEE [IEE93], pages 472–481. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Chatterjee:1993:MRA

[CGS93] Siddhartha Chatterjee, John R.

Gilbert, and Robert Schreiber. Mobile and replicated alignment of arrays in data-parallel programs. In IEEE [IEE93], pages 420–429. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Cheng:1993:HDC

[Che93] Doreen Y. Cheng. Heterogeneous distributed computing — summary of a minisymposium. In IEEE [IEE93], page 762. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Coyne:1993:HPS

[CHW93] Robert A. Coyne, Harry Hulen, and Richard Watson. The high performance storage system. In IEEE [IEE93], pages 83–92. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Choudhary:1993:HPF

[CKZ93] Alok Choudhary, Charles Koelbel, and Mary Zosel. High Performance Fortran: Implementor and users workshop. In IEEE [IEE93], pages 610–614. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Cheng:1993:IRV

[CLF^{+}93] Gang Cheng, Yinguha Lu, Geoffrey Fox, Kim Mills, and Tomasz Haupt. An interactive remote visualization environment for an electromagnetic scattering simulation on a high performance computing system. In IEEE [IEE93], pages 317–326. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Crawfis:1993:VRS

[CMBC93] Roger Crawfis, Nelson Max, Barry Beccker, and Brian Cabral. Volume rendering of 3D scalar and vector fields at LLNL. In IEEE [IEE93], pages 570–576. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Chapman:1993:DDD

[CMMZ93] Barbara Chapman, Piyush Mehrotra, Hans Moritsch, and Hans Zima. Dynamic data distributions in Vienna Fortran. In IEEE [IEE93], pages 284–293. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Clement:1993:APP

[CQ93] Mark J. Clement and Michael J.

Quinn. Analytical performance prediction on multicomputers. In IEEE [IEE93], pages 886–894. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Dagum:1993:APU

[Dag93] Leonardo Dagum. Automatic partitioning of unstructured grids into connected components. In IEEE [IEE93], pages 94–101. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Devine:1993:MPA

[DFWM93] Karen D. Devine, J. E. Flaherty, Stephen F. Wheat, and Arthur B. Maccabe. A massively parallel adaptive finite element method with dynamic load balancing. In IEEE [IEE93], pages 2–11. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM).

ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Douglas:1993:WFP

[DGL^{+}93] David Douglas, Anoop Gupta, Ola Lubeck, David Maier, Paul Messina, Justin Ratner, Burton Smith, and Frederica Darema. What’s in the future for parallel architectures? In IEEE [IEE93], pages 872–873. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Ding:1993:MCS

[Din93] Hong Q. Ding. Monte Carlo simulations of quantum systems on massively parallel supercomputers. In IEEE [IEE93], pages 34–43. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Darnell:1993:CCU

[DK93a] Ervan Darnell and Ken Kennedy. Cache coherence using local knowledge. In IEEE [IEE93], pages 720–729. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Drapeau:1993:SLT

[DK93b] Ann L. Drapeau and Randy H. Katz. Striping in large tape libraries. In IEEE [IEE93], pages 378–387. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Dongarra:1993:LDO

[DPW93] Jack J. Dongarra, Roldan Pozo, and David W. Walker. LA-PACK++: A design overview of object-oriented extensions for high performance linear algebra. In IEEE [IEE93], pages 162–171. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Ewing:1993:DCW

[EMO^{+}93] R. E. Ewing, D. Mitchum, P. O’Leary, R. C. Sharpley, and J. S. Sochacki. Distributed computation of wave propagation models using PVM. In IEEE [IEE93], pages 22–31. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Foster:1993:ITD

[FK93] Ian Foster and Carl Kesselman.

Integrating task and data parallelism. In IEEE [IEE93], pages 154–155. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Fox:1993:CRS

[FRS^{+}93] Geoffrey Fox, Sanjay Ranka, Michael Scott, Allen D. Malony, Jim Browne, Marine C. Chen, Alok Choudhary, Thomas Chetham, Jan Cuny, Rudolf Eigenmann, Amr Fahmy, Ian Foster, Dennis Gannon, Tom Haup, Mike Karr, and Carl Kesselman. Common runtime support for high performance parallel languages parallel compiler runtime consortium. In IEEE [IEE93], pages 752–757. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Galbreath:1993:ADP

[GGL93] N. Galbreath, W. Gropp, and D. Levine. Applications-driven parallel I/O. In IEEE [IEE93], pages 462–471. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Gillilan:1993:SVC

[GL93] Richard E. Gillilan and Bruce R. Land. Scientific visualization of chemical systems. In IEEE [IEE93], pages 296–301. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Graffunder:1993:BPI

[Gra93] Sara K. Graffunder. Barrierbreaking performance for industrial problems on the CRAY C916. In IEEE [IEE93], pages 516–519. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Heller:1993:GBP

[HKS93] D. E. Heller, A. Karp, and H. D.

Simon. Gordon Bell Prize lectures 1993. In IEEE [IEE93], pages 514–515. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Hiranandani:1993:PEF

[HKT93] Seema Hiranandani, Ken Kennedy, and Chau-Wen Tseng. Preliminary experiences with the Fortran D compiler. In IEEE [IEE93], pages 338–350. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Harzallah:1993:HSA

[HS93] Karim Harzallah and Kenneth C.

Sevcik. Hot spot analysis in large scale shared memory multiprocessors. In IEEE [IEE93], pages 895–905. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

IEEE:1993:PSP

[IEE93] IEEE, editor. Proceedings, Supercomputing ’93: Portland, Oregon, November 15–19, 1993. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Mori:1993:DSM

[iMSG^{+}93] Shin ichiro Mori, Hideki Saito, Masahiro Goshima, Mamoru Yanagihara, Takaashi Tanaka, David Fraser, Kazuki Joe, Hiroyuki Nitta, and Shinji Tomita. A distributed shared memory multiprocessor: ASURA — memory and cache architectures. In IEEE [IEE93], pages 740–749. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Jablonowski:1993:VVA

[JBBH93] David J. Jablonowski, John D. Bruner, Brian Bliss, and Robert B. Haber. VASE: The visualization and application steering environment. In IEEE [IEE93], pages 560–569. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM).

ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Karamcheti:1993:CRS

[KC93] Vijay Karamcheti and Andrew Chien. Concert-efficient runtime support for concurrent objectoriented programming languages on stock hardware. In IEEE [IEE93], pages 598–607. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Kaushik:1993:ETA

[KHJ^{+}93] S. D. Kaushik, C.-H. Huang, J. R.

Johnson, R. W. Johnson, and P. Sadayappan. Efficient transposition algorithms for large matrices. In IEEE [IEE93], pages 656–665. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Kahaner:1993:SAW

[KM93] D. X. Kahaner and A. D. Malony. Supercomputing around the world. In IEEE [IEE93], pages 874–876. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Kwan:1993:EGN

[KT93] Thomas T. Kwan and Jeffrey A.

Terstriep. Experiments with a gigabit neuroscience application on the CM-2. In IEEE [IEE93], pages 133–142. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Kwan:1993:CCP

[KTR93] Thomas T. Kwan, Brian K. Toty, and Daniel A. Reed. Communication and computation performance of the CM-5. In IEEE [IEE93], pages 192–201. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Lynch:1993:MPF

[LCL^{+}93] V. E. Lynch, B. A. Carreras, J. N. Leboeuf, B. C. Curtis, and R. L. Troutman. Multi-CPU plasma fluid turbulence calculations on CRAY Y-MP C90. In IEEE [IEE93], pages 308–316. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Li:1993:PES

[LLS^{+}93] Xiquing Li, Gordon Linoff, Stephen J. Smith, Craig Stanfill, and Kurt Thearling. A practical external sort for shared disk MPPs. In IEEE [IEE93], pages 666–675. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM).

ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Long:1993:SBE

[LM93] L. N. Long and J. Myczkowski.

Solving the Boltzmann equation at 61 gigaflops on a 1024-Node CM-5. In IEEE [IEE93], pages 528–534. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Lumetta:1993:DOP

[LML^{+}93] S. Lumetta, L. Murphy, X. Li, D. Culler, and I. Khalil. Decentralized optimal power pricing: The development of a parallel program. In IEEE [IEE93], pages 240–249. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Leutenegger:1993:DCF

[LS93a] Scott T. Leutenegger and Xian-He Sun. Distributed computing feasibility in a non-dedicated homogeneous distributed system. In IEEE [IEE93], pages 143–152. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Liu:1993:SDM

[LS93b] Jie Liu and Vikram A. Saletore.

Self-scheduling on distributedmemory machines. In IEEE [IEE93], pages 814–823. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Lomdahl:1993:GMD

[LTGJB93] Peter S. Lomdahl, Pablo Tamayo, Niels Gronbech-Jensen, and David M. Beazley. 50 GFlops molecular dynamics on the Connection Machine 5. In IEEE [IEE93], pages 520–527. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Lewis:1993:DMM

[Lv93] John G. Lewis and Robert A. van de Geijn. Distributed memory matrix-vector multiplication and conjugate gradient algorithms. In IEEE [IEE93], pages 484–492. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Miles:1993:CIF

[Mil93] Douglas Miles. Compute intensity and the FFT. In IEEE [IEE93], pages 676–684. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Means:1993:BSP

[MWBL93] Robert W. Means, Bret Wallach, David Busby, and Robert C. Lengel, Jr. Bispectrum signal processing on HNC’s SIMD numerical array processor (SNAP). In IEEE [IEE93], pages 535–537. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Machiraju:1993:EFV

[MY93] Raghu K. Machiraju and Roni Yagel. Efficient feed-forward volume rendering techniques for vector and parallel processors. In IEEE [IEE93], pages 699–708. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Norris:1993:SGR

[NP93] Cindy Norris and Lori L. Pollock. A scheduler-sensitive global register allocator. In IEEE [IEE93], pages 804–813. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Naik:1993:PAJ

[NSS93] Viyay K. Naik, Sanjeev K. Setia, and Mark S. Squillante. Performance analysis of job scheduling policies in parallel supercomputing environments. In IEEE [IEE93], pages 824–833. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Netzer:1993:AML

[NX93] Robert H. B. Netzer and Jian Xu. Adaptive message logging for incremental replay of messagepassing programs. In IEEE [IEE93], pages 840–849. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Ni:1993:ISL

[NXK93] Lionel M. Ni, Hong Xu, and Edgar T. Kalns. Issues in scalable library design for massively parallel computers. In IEEE [IEE93], pages 181–190. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Ortega:1993:FDP

[OHA93] Frank A. Ortega, Charles D. Hansen, and James P. Ahrens. Fast data parallel polygon rendering. In IEEE [IEE93], pages 709–718. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Papadopoulos:1993:IBB

[PBGB93] G. M. Papadopoulos, G. A.

Boughton, R. Greiner, and M. J. Beckerle. *T: Integrated building blocks for parallel computing. In IEEE [IEE93], pages 624–635. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Pickett:1993:ESH

[PM93] James K. Pickett and David G.

Meyer. Enhanced superscalar hardware: The schedule table. In IEEE [IEE93], pages 636–644. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Pasquale:1993:SAI

[PP93] Barbara K. Pasquale and George C.

Polyzos. A static analysis of I/O characteristics of scientific applications in a production worklaod. In IEEE [IEE93], pages 388–397. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Ponnusamy:1993:RCT

[PSC93] Ravi Ponnusamy, Joel Saltz, and Alok Choudhary. Runtime compilation techniques for data partitioning and communication schedule reuse. In IEEE [IEE93], pages 361–370. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Pavlakos:1993:VES

[PSM93] Constantine J. Pavlakos, Larry A. Schoof, and John F. Mareda. A visualization environment for supercomputing-based applications in computational mechanics. In IEEE [IEE93], pages 550–559. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Poulsen:1993:ETP

[PY93] David K. Poulsen and Pen-Chung Yew. Execution-driven tools for parallel simulation of parallel architectures and applications. In IEEE [IEE93], pages 860–869. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Ries:1993:PPM

[RAA^{+}93] Bernhard Ries, R. Anderson, W. Auld, D. Breazeal, K. Callaghan, E. Richards, and W. Smith. The Paragon performance monitoring environment. In IEEE [IEE93], pages 850–859. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Rothberg:1993:EBO

[RG93] Edward Rothberg and Anoop Gupta. An efficient blockoriented approach to parallel sparse Cholesky factorization. In IEEE [IEE93], pages 503–512. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Robinson:1993:ECD

[RJMC93] D. F. Robinson, D. Judd, P. K.

McKinley, and B. H. C. Cheng. Efficient collective data distribution in all-port wormholerouted hypercubes. In IEEE [IEE93], pages 792–801. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Ranka:1993:HCH

[RKWF93] S. Ranka, M. Kaddoura, A. Wang, and G. Fox. Heterogeneous computing on heterogeneous systems: Software and application issues. In IEEE [IEE93], pages 763–764. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Raghunath:1993:DIN

[RR93] M. T. Raghunath and Abhiram Ranade. Designing interconnection networks for multi-level packaging. In IEEE [IEE93], pages 772–781. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Stephenson:1993:UVS

[SBS^{+}93] Mary Stephenson, Larry Brandt, Eric Sills, Alex Ropelewski, Bruce Loftis, and Bill Buzbee. Uses of videoconferencing in the supercomputing environment. In IEEE [IEE93], pages 68–71. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Sethian:1993:CFM

[Set93] J. A. Sethian. Computational fluid mechanics and massively parallel processors. In IEEE [IEE93], pages 74–82. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Stathopoulos:1993:RSP

[SF93] Andreas Stathopoulos and Charlotte Froese Fischer. Reducing synchronization on the parallel Davidson method for the large, sparse, eigenvalue problem. In IEEE [IEE93], pages 172–180. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Saavedra:1993:MBA

[SGC93] Rafael H. Saavedra, R. Stockton Gaines, and Michael J. Carlton. Micro benchmark analysis of the KSR1. In IEEE [IEE93], pages 202–213. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Shay:1993:NII

[Sha93] H. D. Shay. National Information Infrastructure (NII) at Supercomputing ’93. In IEEE [IEE93], pages 376–377. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Singh:1993:PAF

[SHHG93] Jaswinder P. Singh, Chris Holt, John L. Hennessy, and Anoop Gupta. A parallel adaptive fast multipole method. In IEEE [IEE93], pages 54–65. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Schwiebert:1993:OFA

[SJ93] Loren Schwiebert and D. N.

Jayasimha. Optimal fully adaptive wormhole routing for meshes.

In IEEE [IEE93], pages 782–791. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Singh:1993:ECK

[SJGH93] Jaswinder Pal Singh, Truman Joe, Anoop Gupta, and John L. Hennessy. An empirical comparison of the Kendall Square Research KSR-1 and Stanford DASH multiprocessors. In IEEE [IEE93], pages 214–225. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Schweitzer:1993:DDM

[SMBZ93] Roland Schweitzer, Jonathan McHugh, Patrick J. Burns, and Charles Zeeb. Daylighting design via Monte Carlo with a corresponding scientific visualization. In IEEE [IEE93], pages 250–259. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Spray:1993:PLR

[SSL^{+}93] Andrew Spray, Heiko Schröder, K. T. Lie, Erik Plesner, and Peter Bray. PIPADS — A low-cost realtime visualization tool. In IEEE [IEE93], pages 398–407. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Stanberry:1993:PCC

[Sta93] Linda Stanberry. Parallel C/ C++: Convergence or divergence. In IEEE [IEE93], pages 448–449. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Singh:1993:PHI

[STB^{+}93] Raj K. Singh, Stephen G. Tell, Shaun J. Bharrat, D. Becker, and V. L. Chi. A programmable HIPPI interface for a graphics supercomputer. In IEEE [IEE93], pages 124–132. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Sabot:1993:PEF

[SWBO93] Gary Sabot, Skef Wholey, Jonas Berlin, and Paul Oppenheimer. Parallel execution of a Fortran 77 weather prediction model. In IEEE [IEE93], pages 538–545. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Tamam:1993:CCC

[TGJ93] Olivier Temam, Elana D. Granston, and William Jalby. To copy or not to copy: A compile-time technique for assessing when data copying should be used to eliminate cache conflicts. In IEEE [IEE93], pages 410–419. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Teng:1993:DAT

[TSBP93] Y. Ansel Teng, Francis Sullivan, Isabel Beichl, and Enrico Puppo. A data-parallel algorithm for three-dimensional Delaunay triangulation and its implementation. In IEEE [IEE93], pages 112–121. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993. VanderWijngaart:1993:EID

[Van93] Rob F. Van der Wijngaart. Efficient implementation of a 3-Dimensional ADI method on the iPSC/860. In IEEE [IEE93], pages 102–111. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Washington:1993:CMG

[Was93] Warren M. Washington. Computer modeling of the global warming effect. In IEEE [IEE93], pages 304–307. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Webb:1993:LBC

[Web93] Jon A. Webb. Latency and bandwidth considerations in parallel robotics image processing. In IEEE [IEE93], pages 230–239. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Wells:1993:QSM

[WOU^{+}93] J. C. Wells, V. E. Oberacker, A. S. Umar, C. Bottcher, M. R. Strayer, J. Drake, and R. Flanery. The Quantum Structure of Matter Grand Challenge Project: Large-scale 3-D solutions in relativistic quantum dynamics. In IEEE [IEE93], pages 44–53. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Warren:1993:PHO

[WS93] Michael S. Warren and John K.

Salmon. A parallel hashed octtree N-body algorithm. In IEEE [IEE93], pages 12–21. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993. URL ftp://

ftp.cacr.caltech.edu/nbody/
sc93.ps.Z.

Zaafrani:1993:PGS

[ZI93] A. Zaafrani and M. R. Ito. Partitioning the global space for distributed memory systems. In IEEE [IEE93], pages 327–336. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.